OpenStudy (anonymous):

Suppose you draw a card from a deck of 52 cards. What is the probability that the card you draw is NOT an ace? There are 4 aces in the deck.

OpenStudy (anonymous):

its the inverse of getting an ace i.e. 1-4/52

OpenStudy (anonymous):

what do i do

OpenStudy (anonymous):

4 in 52 chances of getting an ace. So must be 1- ( 4/52) of not getting an ace. Solving the inverse problem is a useful trick.

OpenStudy (anonymous):

options A. 13/12 B. 12/13 C. 1/13 D. 1

OpenStudy (anonymous):

calculate 1-4/52

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

Which part isnt clear?

OpenStudy (anonymous):

the calculatin part!!!!!!!!!!!!!

OpenStudy (anonymous):

how do i calculate

OpenStudy (anonymous):

1-4/52 = 1-1/13= 12/13

OpenStudy (anonymous):

okayy thanks! (: